# arXiv:2512.22888v2[quant-ph] 28 Feb 2026

### 28 Feb 2026 v2

@@ II Checkerboard code
Since the model is invariant under the exchange $\hat{\sigma}^{x}\leftrightarrow\hat{\sigma}^{z}$, the $\hat{A}_{c}$ and $\hat{B}_{c}$ stabilizers share the same excitation spectrum and dynamics. The model is characterized by a $\mathbb{Z}_{2}$ charge conservation law for each $IJ$-plane of cubes ($IJ=xy,yz,zx$), as indicated by the relation $\prod_{c\in IJ-\text{plane}}\hat{A}_{c}=\prod_{c\in IJ-\text{plane}}\hat{B}_{c}=1$ which enforces that any excitation above the ground state must satisfy these $\mathbb{Z}_{2}$ conservation laws on each plane. Due to the geometry of the checkerboard distribution, the application of an individual $\hat{\sigma}^{x}$ or $\hat{\sigma}^{z}$ operator to a specific qubit generates four elementary excitations, known as fractons, associated with the $\hat{B}_{c}$ or $\hat{A}_{c}$ stabilizers sharing the affected qubit.
These fractons can be viewed as gapped topological defects generated on top of the ground state. Due to overlapping $\mathbb{Z}_{2}$ charge conservation laws, individual fractons cannot be moved without creating additional excitations. However, a pair of fractons (a dipole) possesses restricted mobility in directions orthogonal to its dipole moment. For example, as depicted in Fig. 1(b), an adjacent pair of fractons (a dipole) living in one $xz$-plane can be moved without energy cost along a rigid line l in the $y$-direction by applying string operator $\prod_{n\in\textit{l}}\hat{\sigma}^{x}_{n}$ or $\prod_{n\in\textit{l}}\hat{\sigma}^{x/z}_{n}$. Moreover, if a dipole is oriented along the z-axis, the pair can move freely along the $x$ and $y$ directions. This motion is generated by stacks of adjacent $z$-strings, forming a flexible membrane that permits motion within the $xy$-plane, as shown in Fig. 1(c, d).
@@ III.4 Generalized Duality in classical spin systems
where, $\eta_{j}=-1$ with probability $p$. One can define an averaged Boltzmann weight for each of the cases where the spins of individual coupling $\partial C$ are antialigned for $0\leq k\leq n$ of the $n$ replicas and are aligned in the remaining nkn-k replicas. Averaging over the possible values of $\eta_{C}$, one gets:
@@ IV.1 Random Spin Models With Subsystem Symmetry
Beyond the trivial global $\mathbb{Z}_{2}$ symmetry, the Hamiltonian is invariant under flipping all spins in any arbitrary $XY$, YZYZ, or $XZ$ layer of the lattice. This subextensive set of subsystem symmetries undergoes spontaneous symmetry breaking, leading to a subextensive ground-state degeneracy of $2^{3L-6}$. This degeneracy directly reflects the ability of the checkerboard code to encode 3L63L-6 logical qubits.
@@ IV.3 Numerical Results
We simulated the random Tetrahedral Ising model using large-scale parallel tempering Monte Carlo simulations, averaging the results of enough disorder configurations in order to obtain the disorder averaged features of the model. Each configuration simulated for a given value of the disorder parameter $p$ is obtained by starting from a fully FM set of couplings $(\eta_{\mathbf{v}}^{\pm}=+1\,\,\forall\,\,\mathbf{v})$ and then proposing a random flip to each coupling with probability $p$. This way, the average ratio between number of AFM couplings and the total number of couplings equals $p$.
@@ V Considerations about Haah’s code
For simplicity, we take $\eta^{+}=\eta^{-}=1$, meaning that the two qubit families have the same error probability $p$.

### 03 Mar 2026 v3

# Error Resilience of Fracton Codes and Near Saturation of Code-Capacity Threshold in 3DD

@@ II Checkerboard code
Since the model is invariant under the exchange $\hat{\sigma}^{x}\leftrightarrow\hat{\sigma}^{z}$, the $\hat{A}_{c}$ and $\hat{B}_{c}$ stabilizers share the same excitation spectrum and dynamics. The model is characterized by a $\mathbb{Z}_{2}$ charge conservation law for each IJIJ-plane of cubes ($IJ=xy,yz,zx$), as indicated by the relation $\prod_{c\in IJ-\text{plane}}\hat{A}_{c}=\prod_{c\in IJ-\text{plane}}\hat{B}_{c}=1$ which enforces that any excitation above the ground state must satisfy these $\mathbb{Z}_{2}$ conservation laws on each plane. Due to the geometry of the checkerboard distribution, the application of an individual $\hat{\sigma}^{x}$ or $\hat{\sigma}^{z}$ operator to a specific qubit generates four elementary excitations, known as fractons, associated with the $\hat{B}_{c}$ or $\hat{A}_{c}$ stabilizers sharing the affected qubit.
These fractons can be viewed as gapped topological defects generated on top of the ground state. Due to overlapping $\mathbb{Z}_{2}$ charge conservation laws, individual fractons cannot be moved without creating additional excitations. However, a pair of fractons (a dipole) possesses restricted mobility in directions orthogonal to its dipole moment. For example, as depicted in Fig. 1(b), an adjacent pair of fractons (a dipole) living in one xzxz-plane can be moved without energy cost along a rigid line l in the $y$-direction by applying string operator $\prod_{n\in\textit{l}}\hat{\sigma}^{x}_{n}$ or $\prod_{n\in\textit{l}}\hat{\sigma}^{x/z}_{n}$. Moreover, if a dipole is oriented along the z-axis, the pair can move freely along the $x$ and $y$ directions. This motion is generated by stacks of adjacent $z$-strings, forming a flexible membrane that permits motion within the xyxy-plane, as shown in Fig. 1(c, d).
@@ III.4 Generalized Duality in classical spin systems
where, $\eta_{j}=-1$ with probability pp. One can define an averaged Boltzmann weight for each of the cases where the spins of individual coupling $\partial C$ are antialigned for $0\leq k\leq n$ of the $n$ replicas and are aligned in the remaining nkn-k replicas. Averaging over the possible values of $\eta_{C}$, one gets:
@@ IV.1 Random Spin Models With Subsystem Symmetry
Beyond the trivial global $\mathbb{Z}_{2}$ symmetry, the Hamiltonian is invariant under flipping all spins in any arbitrary XYXY, YZYZ, or XZXZ layer of the lattice. This subextensive set of subsystem symmetries undergoes spontaneous symmetry breaking, leading to a subextensive ground-state degeneracy of $2^{3L-6}$. This degeneracy directly reflects the ability of the checkerboard code to encode 3L63L-6 logical qubits.
@@ IV.3 Numerical Results
We simulated the random Tetrahedral Ising model using large-scale parallel tempering Monte Carlo simulations, averaging the results of enough disorder configurations in order to obtain the disorder averaged features of the model. Each configuration simulated for a given value of the disorder parameter $p$ is obtained by starting from a fully FM set of couplings $(\eta_{\mathbf{v}}^{\pm}=+1\,\,\forall\,\,\mathbf{v})$ and then proposing a random flip to each coupling with probability pp. This way, the average ratio between number of AFM couplings and the total number of couplings equals pp.
@@ V Considerations about Haah’s code
For simplicity, we take $\eta^{+}=\eta^{-}=1$, meaning that the two qubit families have the same error probability pp.